ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 11, 20, 22 Helversen-Pasotto, Anna, CNRS, France Représentations de GL(n, F_{q}) et sommes de Gauss The results of the following two publications will be compared and generalized: A. Helversen-Pasotto, Darstellungen von GL(3, F_{q}) und Gaussche Summen, Math. Ann. 260 (1982) 1- 21 R. W. Carter, Cuspidal Matrix Representations for GL(2, q) and GL(3, q), Proc. London Math. Soc. (3) 64 (1992) 487 - 523